a rectangle has a perimeter of 74 inches. the length of one side is 3X-1.
the width of one side is 2X+3. What is the area?

Answer by keksjr(7)

Area of a Rectangle is A=LW and the Perimeter of a Rectangle is P=2L+2W
Therefore; L=3X-1, and W=2X+3, P=74,
P=2L+2W
74=2(3X-1)+2(2X+3)
74=6X-2+4X+6
74=10X+4
74-4=10X
X=7. finding the value of L and W; 2L=2(3X-1),
2L=2[3(7)-1]
L=20 in.
2W=2(2X+3)
2W=2[2(7)+3]
W=17 in.
Area of the Rectangle A=LW, A = 20x17 = 340sq.in